### v1

# Characterization of the JadePix-3 CMOS pixel sensor for the CEPC vertex detector

Abstract
The Circular Electron-Positron Collider (CEPC), a proposed next-generation $e^{+}e^{-}$ collider to enable high-precision studies of the Higgs boson and potential new physics, imposes rigorous demands on detector technologies, particularly the vertex detector. JadePix-3, a prototype Monolithic Active Pixel Sensor (MAPS) designed for the CEPC vertex detector, targets a spatial resolution of 3 $\rm{\upmu m}$ to meet the key requirement. This paper presents a detailed laboratory-based characterization of the JadePix-3 sensor, focusing on the previously under-explored effects of substrate reverse bias voltage on key performance metrics: charge collection efficiency, average cluster size, and detection efficiency. Systematic testing demonstrated that JadePix-3 operates reliably under reverse bias, exhibiting a reduced input capacitance, an expanded depletion region, enhanced charge collection efficiency, and a lower fake-hit rate. Simultaneously, an extensive methodology for MAPS characterization is established, which determined the optimal operational working conditions for the JadePix-3 sensor. Furthermore, these findings confirm the sensor’s potential for high-precision particle tracking and vertexing at the CEPC while offering valuable references for future iterational R&D of the JadePix series.
keywords:
CEPC , Monolithic active pixel sensors , CMOS , Particle detection
journal: NIM-A
[…]
1 Introduction
The discovery of the Higgs boson has spurred interest in circular $e^{+}e^{-}$ colliders as premier facilities for high-precision studies of the Standard Model Higgs boson and other fundamental particles. Planned designs aim to cover a wide range of center-of-mass energies—from the Z pole (91 GeV) and WW threshold (161 GeV) to the Higgs factory region (240 GeV) and the $t\bar{t}$ threshold (340–365 GeV)—using an $e^{+}e^{-}$ accelerator with a circumference of approximately 100 km. In addition, the same tunnel could later accommodate a hadron collider with a center-of-mass energy nearly 100 TeV. The Circular Electron-Positron Collider (CEPC) [1, 2] has been proposed to pursue this road map, envisioned to investigate the Higgs boson further and reveal new physics potentially.
CEPC’s ambitious physics objectives impose stringent performance requirements on its detectors—particularly the vertex detector. As outlined in Table 1, the CEPC vertex detector must achieve exceptional spatial resolution, maintain a low material budget, minimize power consumption, handle high hit densities, and exhibit robust radiation tolerance to cope with the beam-induced background near the interaction point. In this context, only Monolithic Active Pixel Sensors (MAPS) offer a promising solution to meet these demanding criteria [3, 4].
Therefore, considerable R&D efforts have been undertaken to develop advanced pixel sensor prototypes with monolithic technology for the CEPC vertex detector, such as the TaichuPix series [5, 6, 7] and the JadePix series. The JadePix series is specifically developed to meet the CEPC vertex detector’s demanding criteria, offering high position resolution along with enhanced efficiency. Building on the progressive improvements of its predecessors—JadePix-1, which was to verify the sensor fabrication and optimize the pixel architecture [8, 9], and JadePix-2, which featured an improved analog front‐end design [10]—JadePix-3 represents the latest development and aims for a position resolution of 3 $\rm{\upmu m}$ [11]. The next part presents the detailed design specifications of JadePix-3.
Table 1: Demands for the CEPC vertex detector [1].
Single point resolution $\rm{\upmu m}$
Thickness of sensor chips 50 $\rm{\upmu m}$
Power consumption <50 $\rm mW/cm^{2}$
Timestamp precision $\rm{\upmu s}$
Pixel-hit rate 40 $\rm MHz/cm^{2}$
Radiation tolerance TID, 3.4 Mrad/year
NIEL, $\rm 6.2\times 10^{12}n_{eq}/(cm^{2}year)$
Figure 1: (a) Layout of the entire sensor mask; (b) Circuit topology for the analog front-end [11].
1.1 Overview of JadePix-3 design
The JadePix-3 is designed for a high spatial resolution of 3 $\rm{\upmu m}$. With a pixel matrix consisting of 512 rows and 192 columns, the sensor covers a sensitive area of 8.19 $\times$ 4.85 $\rm{mm^{2}}$ as shown in Fig. 1. The sensor is divided into four distinguishable sectors in the column direction. As presented in Tab. 2, while the row pitch is uniformly set at 16 $\rm{\upmu m}$ across all four sectors, the column pitch varies between 23.11 $\rm{\upmu m}$ in sector 2 and 26 $\rm{\upmu m}$ in sectors 0, 1, and 3. The pixel arrangement in each sector is mirrored, with shared vertical metal lines for biasing and control signals.
Fig. 1 demonstrates JadePix-3’s front-end circuit design. As described in Ref. [11] in detailed, the initial amplification stage utilizes a cascode amplifier, which provides high gain while minimizing noise, with its output referred to as OUT_A. Subsequently, the amplified signal is fed into a discriminator stage, where it is compared against a predefined threshold. The resulting digital output, designated as OUT_D, enables efficient hit detection.
Figure 2: Schematic diagram of the cross-section of the sensor when a reverse bias is applied to the P-substrate via a P-well.
Table 2: Analog front-ends and digital logic blocks for different sectors.
Sector pixel layout digital logic block Remarks
0 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V0 SR flipflop
1 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V1 D flipflop, w/o MASK
2 16 $\times$ 23.11 $\rm{\upmu m^{2}}$ DGT_V2 SR flipflop, w/o MASK
3 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V0 Increased current for the front-end
The chip employs both rolling shutter (RS) and global shutter (GS) readout sequence modes. In the conventional rolling shutter mode, one row is activated for reading at a time [12]. This process requires 192 $\rm ns$ to transmit hit data and reset the pixels, and scanning the entire 512-row matrix takes 98.3 $\rm{\upmu s}$. While in global shutter, all 512 rows are frozen for a period of time during readout.
The JadePix-3’s sensor design is shown in Fig. 2. A reverse bias $V_{\rm sub}$ can be applied on P-well and P-substrate, expanding the depletion region around the N-well electrode.
1.2 Characterization Motivation
Previous initial characterization proved that JadePix-3 had achieved the general design requirements under zero substrate reverse bias. The result revealed minimum thresholds ranging from 90 to 140 electrons, with a fake-hit rate below $\rm 1\times 10^{-10}$ per frame per pixel [11]. In the beam test for the JadePix-3 telescope, the sensor showed spatial resolutions of 5.2 $\rm{\upmu m}$ and 4.6 $\rm{\upmu m}$ in two dimensions, with a maximum detection efficiency approaching 99.0%  [13]. However, the full performance potential of the JadePix-3 sensor remains incompletely characterized, particularly concerning the effects of reverse bias.
The performance of MAPS is significantly influenced by the bias voltage applied to the substrate. This bias has a multifaceted impact: on one hand, it directly increases the size of the depletion zone and reduces the sensor’s input capacitance [14, 15]. On the other hand, it also affects the characteristics of the integrated circuitry; specifically, as the substrate bias becomes more pronounced, the threshold of NMOS transistors within the pixel circuitry tends to increase due to the body effect, consequently altering the circuit’s operating status. Therefore, adjusting the circuit’s operating point after applying substrate bias and investigating the resulting impact on these key performance metrics becomes essential. These metrics are vital for the R&D of the CEPC vertex detector.
Consequently, this work provides a thorough study of reverse bias effects, which is imperative for optimizing sensor design and ensuring its suitability for high-precision particle tracking applications.
This work also introduces an integrated and systematic test methodology (detailed in Section 2), which was established to guide the subsequent comprehensive, lab-based characterization of MAPS behavior under varying reverse bias conditions. While most studies hold the substrate bias fixed or explore a narrow range [16, 17], our approach systematically sweeps the reverse bias from 0 V to –6 V under identical threshold settings. This methodology establishes comparable operating conditions across different bias voltages, thereby facilitating an unambiguous assessment of performance changes predominantly attributable to the bias variation itself.
2 Workflow for Comprehensive Characterization
This work employs a structured, integrated methodology (see Fig. 3):
  • 1.
    Phase 1: Measurement Setup, which is introduced in Section 3, focuses on the preparation stage and establishes a stable experimental setup that minimizes mechanical and environmental perturbations, thereby ensuring reliable test conditions.
  • 2.
    Phase 2: Working-Point Tuning is introduced in Section 4. This phase establishes a series of optimal operating points through a two-stage procedure.
    1. (a)
      Target threshold: Target working-points are defined as distinct operating configurations, one for each applied bias level, that all share a consistent average threshold. This is achieved by applying reverse bias and meticulously tuning DAC parameters to meet the predefined target threshold.
    2. (b)
      ITHR scan: Subsequently, an ITHR scan is performed from each of these target working-points to generate an expanded set of conditions, offering a range of distinct threshold levels for each bias.
    This tuning not only confirms the feasibility of operating under varying reverse biases but also lays the groundwork for all subsequent measurements.
  • 3.
    In Phase 3: Comprehensive Performance Evaluation, which is illustrated in Section 5, three sequentially interlinked investigations were carefully designed and performed:
    1. (a)
      Sensor Electrical Properties Characterization: Detailed measurements of leakage current and input capacitance—both regarded as critical sensor performance metrics—were conducted to understand the intrinsic electrical behavior of the sensor under bias.
    2. (b)
      Noise Performance Analysis: Fake-hit rate is measured to assess noise characteristics, since maintaining a low fake-hit rate is critical for ensuring the effectiveness of hit data for the subsequent measurements.
    3. (c)
      Detection Response Evaluation: The sensor’s response is further assessed by exposing it to external sources, including radioactive sources and an infrared laser. The laser tests, in particular, provide high-resolution mapping of in-pixel charge collection efficiency and serve as a proxy for evaluating sensor performance in response to external stimuli.
Figure 3: Detailed workflow to comprehensively characterize the MAPS sensor.
This integrated workflow not only delivers a robust, reproducible characterization of JadePix-3 in the laboratory—fully capturing how reverse bias shapes every aspect of its response—but also provides a transferable framework for future MAPS R&D.
3 Experimental Setup
The JadePix-3 chip has been wire-bonded to an adaptor board and then connected to an external FPGA-based DAQ system (see Fig. 4). The FPGA board configures the chip parameters and transfers data to the PC via Ethernet. A shock-absorbing platform was employed to maintain system stability and prevent displacement of the test source, either a radioactive source or a laser source, and the sensor during the measurements. A source meter is used to apply reverse bias ranging from 0 to -6 V to the substrate.
Figure 4: (a) Schematic view of the test system; (b) Setup of the infrared laser test system. All devices are fixed or placed on an optical table.
To proceed laser experiment under the demand of high spatial resolution, a laser test system featuring three high-precision motor controller is deployed (see Fig. 4). The spatial position in three orthogonal directions of the laser source is regulated precisely by these three motors, each has a repeatable position resolution of 1 $\rm{\upmu m}$ and a minimum step size of 0.1 $\rm{\upmu m}$. Details about the experiment parameters and test method will be discussed thoroughly in Section 5.5.
4 Working-point Selection
This section details the methodology for selecting operational working-points of the JadePix-3 sensor. The primary objective is to establish uniform thresholds that enable the comparison of the sensor’s performance across varying substrate reverse biases. Central to this process is the accurate measurement of pixel thresholds, achieved using the S-curve method.
This methodology involves giving each pixel a certain number of electrical pulses and counting the hits while increasing the injected charge, which is regulated as $V_{\rm pulse}\cdot C_{\rm test}$, where $C_{\rm test}$ is the injection capacitor that has a designed value of $0.18~{}{\rm fF}$ marked in Fig. 1, and $V_{\rm pulse}$ is the external voltage for test-pulse injection.
As the injected charge approaches the pixel threshold, the hit count rises along an S-shaped curve, typically modeled by fitting an error function. The pixel threshold is then defined as the injected charge level at which this fitted error function reaches half of its maximum value. Threshold variations between pixels are quantified by the fixed pattern noise (FPN), defined as the standard deviation of the mean threshold. Furthermore, the temporal noise (TN) is characterized by the RMS value of the derivative of the S-curve, providing a measure of the sum of both the fluctuations on the injected signal and the fluctuations on the pixel threshold itself. The final threshold value is converted to electrons via the designed value of $C_{\rm test}$.
4.1 Target Working-Point Tuning
To thoroughly assess the impact of reverse bias on sensor performance, it is essential to tune the sensor to target working-points that exhibit a uniform average threshold across various substrate biases. Achieving such consistency poses a challenge due to the complex interplay of multiple parameters influencing the analog front-end. Therefore, the operating conditions—particularly the DAC settings that control the front-end state—must be meticulously adjusted.
This work focuses on the characterization of sectors 0-2. Sector 3 is excluded from the present discussion; its front-end current was increased for radiation tolerance evaluations, leading to significantly higher power consumption. The analog front-end design is identical for sectors 0-2, with variations existing only in their digital logic blocks for pixel functions and in their respective pixel pitches (see Tab. 2).
The M6 MOSFET, biased by VCLIP, clamps excessively high OUT_A signals (see Fig. 1). Given that the amplitude of OUT_A increases significantly with reverse bias due to a reduction in sensor capacitance, VCLIP is set to 1.307 V. This is essentially its maximum value and is chosen to eliminate the clamping effect of M6 on the OUT_A amplitude during characterization. To compensate for the increase of the M8 gate threshold induced by substrate reverse bias, VCASN2 is increased from its nominal 400 mV to 722 mV, thereby restoring M8’s switching threshold to its intended value.
[…]
Table 3: Nominal and Target operating configurations for the analog front-end. The seven Target working-points share constant ITHR, VCLIP, and VCASN2 settings, while VCASN is adjusted to achieve a similar average threshold across different $V_{\rm sub}$ levels.
Settings $V_{\rm sub}$ VCASN/mV ITHR/nA VCLIP/mV VCASN2/mV
Nominal 0 V 400 500 0 400
Target 0 V 400 500 1307 722
-1 V 592
-2 V 732
-3 V 861
-4 V 986
-5 V 1129
-6 V 1265
The threshold distributions for these target working-points are presented in Fig. 5. It is observed that these distributions at non-zero reverse biases are noticeably broader compared to zero bias, indicating increased pixel-to-pixel threshold variations. This observation is further substantiated by the FPN analysis shown in Fig. 6, where FPN increases with reverse bias and quantitatively reflects the growing differences in threshold across the pixel matrix. The corresponding TN analysis is shown in Fig. 6. An increase in TN is observed under reverse bias conditions. The underlying reasons for this behavior will be discussed in Section 5.2 along with the change of OUT_A signal amplitude.
[…]
Figure 6: (a) Average FPN as a function of reverse bias. (b) Average TN as a function of reverse bias.
4.2 ITHR Scan
To investigate sensor performance across a range of detection thresholds, the ITHR value was systematically scanned starting from each of these target working-points. This procedure yielded a series of operating conditions—7 distinct $V_{\rm sub}$ levels, each with 11 different ITHR settings—resulting in a set of working-points with varying thresholds.
Figure 7 illustrates the impact of the ITHR value on the mean threshold, FPN and TN, respectively. Both FPN and TN generally exhibit an increasing trend with higher threshold values.
The linear relationship between the mean threshold and the ITHR setting, depicted in Fig. 7, is consistently observed across all tested reverse bias voltages. This linearity not only confirms the consistency of the threshold tuning but also underscores that ITHR is a suitable parameter for threshold control under different reverse biases.
Figure 7: ITHR scan results. (a) Threshold, (b) FPN (fixed pattern noise), and (c) TN (temporal noise) versus ITHR value for working-points of distinct reverse biases.
5 Experimental Results and Discussion
Reducing sensor capacitance and expanding the depletion zone are the most intrinsic and direct consequences of applying a reverse bias on the substrate. After adjusting thresholds under different reverse biases and determining a set of working-points by ITHR value scan described in Sec. 4.2, a thorough sensor characterization regarding the impact of the substrate reverse bias voltage was conducted, including leakage current, analog response, fake-hit rate and charge collection efficiency.
5.1 Leakage Current
In the absence of ionizing radiation, thermally generated free charge carriers are collected by the diode, resulting in what is known as leakage current or dark current. Leakage current from the pixel area impacts the sensor’s electrical performance by inducing shot noise [18].
[…]
5.2 Input Capacitance
The effective pixel input capacitance of the sensor, $C_{\rm effective}$, can be determined by analyzing the analog response of the sensor under different reverse bias voltages. The analog output amplitude (denoted as OUT_A in Fig. 1) is given by:
$A_{\rm{OUT\_A}}=\frac{C_{\rm test}\cdot V_{\rm pulse}}{C_{\rm effective}}\cdot f_{\rm gain}\cdot f_{\rm sf}~{},$ (1)
[…]
Two independent and complementary methods were employed in this study to measure $C_{\rm effective}$ via the relation:
$C_{\rm effective}=C_{\rm test}\cdot\frac{V_{\rm pulse}}{A_{\rm{OUT\_A}}}\cdot f_{\rm gain}\cdot f_{\rm sf}~{}.$ (2)
  • 1.
    Method 1: In this approach, the pulse voltage was fixed at $V_{\rm pulse}=160{\rm mV}$ while the analog response amplitude was recorded via an oscilloscope. The measurement was repeated to extract a statistically robust value for the amplitude. The measured amplitude is then substituted into Eq. (2), the term $A_{\rm{OUT\_A}}$, to obtain $C_{\rm effective}$.
  • 2.
    Method 2: In this procedure, the analog response amplitude, $A_{\rm{OUT\_A}}$, was maintained at a constant value by dynamically adjusting $V_{\rm pulse}$. The corresponding waveform measurement was repeated, and the adjusted $V_{\rm pulse}$ value (required to hold the amplitude constant) was then used in Eq. (2) to compute $C_{\rm effective}$.
Figure 9: Effective capacitance measured as a function of the reverse bias voltage using two methods. (a)Method 1: $V_{\rm pulse}$ is fixed at 160 mV, and the analog response amplitude is measured over 200 repetitions. (b) and (c)Method 2: The analog amplitude is fixed at 49 mV, 78 mV, respectively, by adjusting $V_{\rm pulse}$; measurements are performed at two different current threshold settings (500 pA and 650 pA, respectively). For both methods, the calculation of $C_{\rm effective}$ assumes $f_{\rm gain}=20$ and $f_{\rm sf}=0.7$.
The two methods provide complementary insights. While Method 1 directly captures the amplitude response at a fixed pulse voltage, Method 2, by contrast, keeps the amplitude-dependent $f_{\rm gain}$ factor via maintaining the amplitude value constant, which offers a more reliable $C_{\rm effective}$ result with increasing reverse bias voltage. To examine how $f_{\rm gain}$ varies with amplitude, Method 2 was repeated at two different output levels (49 mV and 78 mV), shown in Fig. 9 and Fig. 9. The results show consistent trends with only tiny differences between the two amplitudes, confirming the reliability of the approach.
Results for these two methods after calculation are shown in Fig. 9. The input capacitance can be reduced by more than 50% when going from $V_{\rm sub}=0~{}{\rm V}$ to $V_{\rm sub}=-2~{}{\rm V}$, corresponding to the rapid expansion of the depletion region. This reduction is very important as smaller input capacitance increases the charge-voltage conversion coefficient at the input of the front-end circuit for a given charge. To keep the input-referred threshold constant, the VCASN was tuned to lower the OUT_A voltage baseline as the OUT_A amplitude increased with reverse bias voltage. This may be related to the increase of FPN and TN for reverse bias from -1 V to -5 V in Fig. 6 6 respectively, as the same effect had been observed in a VCASN scan shown in Ref. [11] (Fig. 13).
5.3 Fake-hit Rate Test
Fake-hit, i.e., spurious signals that do not correspond to any external particle interactions, are a critical parameter in pixel detector performance. These false signals often arise from electronic noise, crosstalk, or other sensor imperfections, leading to erroneous track reconstruction. Therefore, assessing and reducing the fake-hit rate is crucial for ensuring the reliability of tracking detectors.
To compare two distinguished readout sequence modes, the fake-hit rate for each sector $i$ in both rolling shutter (RS) and global shutter (GS) modes is calculated using a unified formula:
$R_{i}^{\text{mode}}=\frac{N_{i}^{\text{mode}}}{N_{\text{frames}}^{\text{mode}}\cdot N_{\text{pixels}}},$ (3)
where $N_{i}^{\text{mode}}$ represents the total number of false hits in sector $i$, $N_{\text{frames}}^{\text{mode}}$ is the number of frames collected, and $N_{\text{pixels}}=512\times 48$ is the number of pixels per sector. In the rolling shutter mode (RS), 100,000 frames are collected ($N_{\text{frames}}^{\text{RS}}=100,000$) in the absence of external sources. In the global shutter mode (GS), 1,000 frames are collected ($N_{\text{frames}}^{\text{GS}}=1,000$) with an acquisition window of 98.3 $\upmu\rm s$, matching the rolling shutter readout time, to enable a comparable measurement.
The sensitivity limit is defined as the lowest measurable fake-hit in the measurement for these two modes and is selected as $\rm 1\times 10^{-10}$ for the rolling shutter and $\rm 1\times 10^{-8}$ for the global shutter mode.
As the fake-hit rate is expected to be strongly influenced by the threshold, a threshold scan under different reverse biases was performed. The results are shown in Fig. 10. In Fig. 10, up to 10 pixels (accounting for 0.04%) in each sector with the highest fake-hit counts are identified as hot pixels that contribute most to the false hits [20, 21]. These hot pixels are masked. In comparison, values presented in Fig. 10 are obtained without masking any pixels.
According to these results, the difference between the two modes is insignificant above the GS sensitivity limit. Reverse bias substantially reduces the fake-hit rate compared with the unbiased state, with $V_{\rm sub}=-3~{}{\rm V},-4~{}{\rm V}$ yielding the lowest values.
Figure 10: Fake-hit Rate measured with RS (rolling shutter) and GS (global shutter) modes versus threshold under varied reverse biases with all pixels included in Fig. 10 and with hot pixels masked in Fig. 10. In Fig. 10, up to 10 pixels (accounting for 0.04%) for each sector, which have the largest counts of fake-hit, are considered hot pixels and masked. Sensitivity limits are indicated as dashed horizontal lines in every pad.
5.4 Radioactive Source Test
Two radioactive sources, ${}^{90}\rm Sr$ and ${}^{55}\rm Fe$, were employed to assess the sensor’s response to external ionizing radiation. ${}^{90}\rm Sr$, a $\beta$ emitter, generates minimum ionizing particles that traverse the sensor and deposit charge over an extended region. This distributed ionization is particularly sensitive to changes in the sensor’s depletion characteristics and charge-sharing behavior under reverse bias, thereby providing insights into the sensor’s charge collection efficiency. In contrast, ${}^{55}\rm Fe$ emits 5.9 keV X-rays that interact with silicon primarily via the photoelectric effect and deposit an ionization charge of 1640 $e^{-}$ in a small point-like spatial region compared to the dimensions of the sensor, resulting in highly localized charge deposition. This characteristic makes ${}^{55}\rm Fe$ an ideal probe for evaluating the sensor’s response to concentrated ionization.
5.4.1 Test with ${}^{90}\rm Sr$ Source
A ${}^{90}\rm Sr$ source was placed 2 mm above the sensor to verify JadePix-3’s response to beta rays. All data were collected with a 100,000-frame rolling shutter scan, ensuring equivalent collection time for every tested working-points. Approximately 30,000 clusters were recorded in sector 0,1,2 over the course of 50 seconds, for which the hit map is shown in Fig. 11. Fig. 12 displays the data analysis result. Due to varying solid angles relative to the fixed source, each sector inherently accumulates a different number of hits. However, because each working-point was recorded with the same number of frames (and thus identical detection time), and radioactive source was kept stationary throughout the measurements, the accumulated cluster count within a given sector is directly proportional to its detection efficiency. Therefore, for each sector, the relative detection efficiency at a given bias voltage was obtained by normalizing its accumulated cluster count by the maximum cluster count recorded for that same sector across all bias settings.
Figure 11: Hit map of sector 0,1,2 from the ${}^{90}\rm Sr$ radioactive source test. Roughly 30,000 events were recorded under $\beta$-ray irradiation from the backside during a time duration of 50 s, with ITHR value fixed at 500 pA and $V_{\rm sub}$ at 0 V.
Figure 12: ${}^{90}\rm Sr$ radioactive source response result. (a) Average cluster size versus the threshold. (b)-(d) Relative detection efficiency versus the threshold for Sector 0, 1, 2, respectively. The relative detection efficiency is proportional to the accumulated number of clusters detected and normalized by the maximum number of clusters measured in each sector.
Fig. 12 demonstrates a clear reduction in cluster size with the increase of the bias voltage, e.g., from 16-17 to 8-9 at 100 $e^{-}$ threshold from 0 V to -6 V. This means more charge from particle interactions is collected within the seed pixel, reducing the spread to neighboring pixels. Additionally, the stronger electric field lowers lateral diffusion, causing the charge to be collected more locally. As a result, fewer pixels surrounding the seed pixel exceed the detection threshold, leading to smaller average cluster size. The slope of cluster size as a function of the threshold also decreases significantly in Fig. 12. This phenomenon demonstrates that the charge collection ability of the sensor is less affected by the threshold under biased conditions, which further illustrates that the charge collection is more sufficient and robust.
Fig. 12-12 demonstrate an enhancement in detection efficiency as reverse bias increases from 0 V to -5 V. This behavior can be attributed to the depletion zone expansion. However, the measured results also exhibit an evident efficiency drop as the bias is further increased from –5 V to –6 V. One possible reason for this decline is that under excessive reverse bias, the characteristics of the MOSFETs may undergo significant changes due to the body effect, leading to a reduction in signal amplification performance. A similar effect under high bulk bias was also observed in Ref. [22], where an nMOS device biased at –6 V exhibited a 260 mV threshold voltage shift and a 40% drop in drain current. This suggests that the analog front-end performance of JadePix-3 may be compromised under such high bias conditions, highlighting the importance of carefully selecting the operational bias range.
5.4.2 Test with ${}^{55}\rm Fe$ Source
A ${}^{55}\rm Fe$ source was positioned 8 mm above the sensor to verify JadePix-3’s response to X-ray. The data collection and processing for ${}^{55}\rm Fe$ were identical to ${}^{90}\rm Sr$, using a 100,000-frame rolling shutter scan, recording approximately 30,000 clusters over 50 seconds, ensuring a consistent collection time window.
Figure 13: ${}^{55}\rm Fe$ radioactive source response result. (a) Average cluster size versus the threshold. (b)-(d) Relative detection efficiency versus the threshold of Sector 0, 1, 2, respectively. The relative detection efficiency is obtained in the same method as for the ${}^{90}\rm Sr$ source test as shown in Fig. 12.
As depicted in Fig. 13, the slope of the cluster size versus threshold curve decreases significantly when $V_{\rm sub}$ increases, providing similar results to the ${}^{90}\rm Sr$ source test, which further illustrates the charge collection is more sufficient under the biased condition. Fig. 13-13 shows improvement in detection efficiency as substrate reverse bias increases.
Comparing Fig. 13-13 internally, sector 1 exhibits a smaller efficiency drop with increasing the threshold (approximately 10%) compared to sectors 0 (20%) and 2 (25%). A similar behavior is also observed for the ${}^{90}\rm Sr$ source test, as shown in Fig. 121212. A plausible mechanism involves the difference of the OUT_D trigger design: sector 1 uses a D-flipflop, while both sector 0 and 2 use SR-flipflops (see Tab. 2). D-flipflop is edge-triggered. While the SR-flipflop is level-triggered or latched. The discrepancy may indicate that the electrical test pulses are different from the ionization signal because the thresholds are calibrated by the electrical pulse while they showed different results on source measurements.
5.5 Infrared Laser Test
Following the configuration of the laser test platform outlined in Section 3, a systematic investigation of the in-pixel detection efficiency was performed using an infrared laser beam. The laser operates at a wavelength of 1064 nm. The laser’s pulse energy is adjustable between 0 and 60 pJ, modulated via an attenuation factor ranging from 0% tune (maximum pulse energy) to 100% tune (minimum pulse energy), controlled through the Laser Power Tune function of the system. For this laser measurement, the selected pixel is located in sector 2, which has the smallest pixel pitch among the sectors (see Tab. 2).
5.5.1 Beam Position Adjustment
To ensure optimal focusing and achieve the minimum laser spot size, which can be tuned to 3.4 $\rm{\upmu m}$, significantly smaller than the pixel size [11], at the depth of the chip’s sensitive layer, a scan was performed along the Z-direction (perpendicular to the sensor plane) to locate the beam waist. This scan involves measuring the number of pixel hits as a function of the laser’s Z-position, which indicates the vertical distance to the sensor surface, at two attenuation levels: 5% and 50%. As shown in Fig. 14, the minimum number of hits, indicative of the beam waist, was observed at a Z-coordinate of 11.65 mm. This position was subsequently fixed for all subsequent measurements to ensure optimal focus and minimal spot size on the chip.
Figure 14: Number of pixel hits as a function of the laser position Z. The upper and lower curves are for the laser power attenuation at 5% and 50%, respectively. Settings of this experiment: laser frequency = 25 kHz, sensor threshold = 215 $e^{-}$, $V_{\rm sub}=0{\rm V}$, Rolling Shutter frame numbers = 1,000.
To approach the center of a single pixel, a two-dimensional XY-scan along the sensor surface was performed, as illustrated in Fig. 15. The cluster size during this scan was restricted to either 0 or 1 under a 75% laser tune setting, ensuring precise delineation of the pixel boundaries. The experimental conditions included a laser frequency of 25 kHz, a threshold of 215 $e^{-}$ for the sensor, a substrate bias voltage $V_{\rm sub}$ of 0 V, and a total of 10,000 Rolling Shutter frames. This scan confirmed the exact center position of a single pixel in both the X- and Y-directions.
Figure 15: XY-scan results from a laser source. The plot on the left shows the hit map, the number of laser pulse events recorded at a specific (X, Y) position. The upper right panel displays the hit profile along the line at Y = 9.774 mm, while the lower right panel shows the hit profile along the line at X = 20.263 mm. Settings of this experiment: laser frequency = 25 kHz, sensor threshold = 215 $e^{-}$, $V_{\rm sub}=0{\rm V}$, Rolling Shutter frames = 10,000.
5.5.2 In-Pixel Detection Efficiency
Building on previous beam test results [13], we further probed the in-pixel detection performance by using the attenuation-tunable infrared laser as an accessible proxy for detection efficiency and charge collection efficiency in response to external excitation.
Based on the XY-scan results as shown in Fig. 15, four distinct in-pixel positions were selected and marked: (A) the pixel center, (B) the center of the long edge, (C) the center of the short edge, and (D) the diagonal center of $2\times 2$ neighbouring pixels. These four positions, annotated in Fig. 15, were representative of distinct in-pixel locations and were selected to investigate spatial variations in detection efficiency.
As the infrared laser penetrates the silicon bulk and generates electrons and holes along the way, we assessed the pixel’s detection efficiency in response to it to mimic the pathways of charged particles. An RS scan was performed over 100,000 frames (40 seconds) under a 25 Hz laser operation for each reverse bias value across the four in-pixel positions. Given that 1000 laser pulse triggers are expected during this period, the detection efficiency is calculated as the ratio of the measured hit count to 1000, thereby quantifying the pixel’s ability to detect individual laser pulses. Results shown in Fig. 16 demonstrate that applying a reverse bias from 0 to –2 V had already enhanced the detection efficiency at in-pixel positions A, B, and C to nearly 100%, underscoring the performance improvement enabled by the reverse bias. These observations are in agreement with beam tests of other MAPS sensors [23], providing an imperative reference to further beam test results.
Figure 16: Detection efficiency in response to laser beam as a function of the reverse bias $V_{\rm sub}$. Measured at four in-pixel spots A, B, C, and D as marked in Fig. 15.
Furthermore, a laser with tuning method was developed and performed in order to quantitatively characterize the sensor’s charge collection capability. The most straightforward way would be to measure the charge deposit in the sensor under external excitation, however, JadePix-3 employs a binary digital readout, and it lacks the analog output of the collected charge. Therefore, an external probe was required whose intensity can be continuously varied to infer the pixel’s charge collection sensitivity. By using the laser’s tunable attenuation as a proxy for the amount of injected charge, the sensor’s binary readout can be translated into a quantitative efficiency index in a method described as follows. The laser frequency was set to 25 Hz, and the chip was operated with 20,000 Rolling Shutter frames, equivalent to an 8-second acquisition window. Under these conditions, if every laser pulse generates enough charge that exceeds the detection threshold, 200 hits would be recorded by the pixel. The laser pulse energy was then attenuated by incrementally increasing the tune value until the number of recorded hits dropped to around 100333The actual experimental values are not exactly 100 and range between 90 and 110 due to unavoidable fluctuation.. The tune value at this point, designated as the Detection Sensitivity Index (DSI), was recorded for each of the four marked positions. This DSI is defined as the laser tune value at which the number of detected hits drops to approximately 100 out of 200 possible triggers; a higher DSI indicates that the pixel remains responsive even at lower laser intensities, thus signifying more effective charge collection efficiency and thus improved detection performance.
Figure 17: Detection Sensitivity Index (DSI) as a function of the reverse bias $V_{\rm sub}$. Measured at four in-pixel spots A, B, C, and D as marked in Fig. 15.
The variation of the DSI with reverse bias $V_{\rm sub}$ was measured and is depicted in Fig. 17. The resulting curves reveal distinct trends across the four in-pixel positions, providing more intricate information about the in-pixel variance. Position A, the pixel center, consistently exhibited the highest DSI, indicating peak detection efficiency, whereas Position D showed the lowest DSI, reflecting diminished sensitivity at the pixel boundary. Position B (long-edge center) and C (short-edge center) displayed intermediate DSI, with Position B substantially outperforming Position C. This aligns with findings from the JadePix-3 beam test measurement, which revealed notably higher detection efficiency at the long-edge center (Position B) compared to the short-edge center (Position C), where efficiency was significantly lower and closely resembled that of the pixel corner (Position D). Notably, the DSI increased with increasing reverse bias across all positions, suggesting enhanced charge collection efficiency at higher $V_{\rm sub}$ due to the expanding depletion volume.
5.6 Optimal Reverse Bias Setting
A comprehensive survey of the JadePix-3’s performance over substrate biases from 0 to –6 V indicates that a reverse bias of approximately –5 V offers the best overall trade-off. This optimal working-point not only improves detection efficiency and charge collection ability, but also maintains a low fake-hit rate and input capacitance.
6 Conclusions and Outlook
To further characterize the performance of JadePix-3, a prototype MAPS developed for the CEPC vertex detector, detailed laboratory-based measurements were performed, in particular with substrate reverse bias voltages applied. A comprehensive test workflow was developed and implemented for the systematic characterization of JadePix-3. Experimental results show that the JadePix-3 sensor can work stably when a substrate reverse bias is applied. Moreover, the reverse bias has a considerable influence on the performance of the sensor, typically being beneficial. Our results reveal that applying reverse bias substantially enhances sensor performance, reducing input capacitance, enlarging the depletion region, and improving charge collection efficiency, while also lowering fake-hit rates. Optimal reverse bias setting is also determined within this integrated test procedure. This test methodology not only confirms JadePix-3’s potential for further performance improvement but also establishes a robust methodology for the evaluation and optimization of similar MAPS devices in future particle tracking applications.
This result acquired in laboratory conditions will also provide critical insights to inform subsequent beam test configurations, parameter optimization, as well as the radiation hardness test. Moreover, this work will serve as the architectural and procedural foundation for the tuning, and optimization of the next iteration, the JadePix-5 sensor, thereby driving forward MAPS development to meet CEPC’s requirements.
Acknowledgement
This work was supported in part by the National Key Research and Development Program of China under Grant 2022YFA1602203 and Grant 2023YFA1606300, the Fundamental Research Funds for the Central Universities of China under Grant WK2360000014, and the STCF Key Technology Research and Development Project. We thank the Hefei Comprehensive National Science Center for its strong support on the promotion and development of the STCF project.